Circular No. 8111

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003dc, 2003dd, 2003de Further to IAUC 8106, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and S. Brar report apparent supernovae found on unfiltered NEAT images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003dc Mar. 26.46 13^{h}54^{m}21.^{s}70 + 5^{o}46′09′′.3 19.3 1′′.2 E, 1′′.5 S
2003dd Mar. 26.49 13 56 55.04 + 3 55 58.2 20.1 14′′ E, 3′′ S
2003de Apr. 1.14 10 10 41.25 +35 44 52.9 17.2 1′′.7 W, 4′′.5 N

Additional magnitudes from NEAT images: SN 2003dc, 2002 May 6.28, [20.9; 2003 Feb. 6.65, [20.5; Mar. 24.58, 19.6; Apr. 3.47, 18.3. SN 2003dd, 2002 June 24.21, [20.9; 2003 Feb. 6.66, [20.1; Apr. 3.46, 19.8. SN 2003de, 2002 Mar. 30.25, [21.1; 2003 Mar. 3.36, [21.0; 21.33, [20.6; Apr. 7.17, 17.1.

SUPERNOVAE 2003cr, 2003da, AND 2003db R. J. Foley, M. Papenkova, D. Weisz, A. V. Filippenko, and R. Chor-

nock, University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Apr. 8 UT with the Shane 3m telescope at Lick Observatory, reveals that SN 2003cr (IAUC 8103) is probably of type Ic, within a few weeks past maximum brightness. SN 2003da (IAUC 8108) is of type II shortly after the explosion; the very blue continuum is nearly featureless, but low-contrast hydrogen Balmer lines and weak He I 587.6-nm are visible and have P-Cyg profiles. SN 2003db (IAUC 8108) is also of type II, probably a few weeks after the explosion. The broad Hα emission line is much stronger than the absorption component.

M. Schwartz and P. R. Holvorcem report the independent LOTOSS (cf. IAUC 7906) discovery of SN 2003db on unfiltered Tenagra II CCD images taken on Apr. 7.4 (at mag 18.1) and 8.2 UT (mag 17.9); the new object was not present (to mag ∼ 18.5) on an image taken on Feb. 28.5.

COMET C/2002 W8 (SOHO) Further to IAUC 8106, D. Hammer reports measurements for a Kreutz

sungrazing comet found on C2 and C3 SOHO website images by K. Černis (not X.-M. Zhou, as stated on MPEC 2003-G32, where the reduced astrometry and orbital elements appear).

2002 UT  α_{2000}  δ_{2000}
Nov. 22.346  15^{h}48.^{m}3  −23^{o}35′
2003 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green